The DVs are x1 and x2, and the OF is -54x2-3x132 +22. What is the Hessian?

Where DV is Desired Value and OF is Objective Function
Answer #1


Hessian iS χχ So gzi.= 5-3x2, fx2ー4-3x1+4x2 Honce Hessian s =10-3 1 3 O-S